### v5

@@ 5.5. Agriculture
GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation (Kamilaris and Prenafeta-Boldú, 2018; Costa et al., 2023). Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage (Sapkota et al., 2024; Kamilaris and Prenafeta-Boldú, 2018), crop yield prediction models that integrate weather, soil, and satellite data (Chlingaryan et al., 2018; Ma et al., 2023), pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy (Mohanty et al., 2016; Li et al., 2022), virtual agronomy advisors that deliver context-sensitive recommendations to farmers (Purdy and Gupta, 2023; Mahdavinejad et al., 2018), agricultural robotics for automated harvesting and monitoring (Duckett et al., 2018; Bechar and Vigneault, 2021), crop breeding simulations that accelerate trait discovery (Varshney et al., 2021; Beltran et al., 2022), and weather forecasting for climate adaptation (Rasp et al., 2020; Nguyen et al., 2023). Recent deployments provide evidence of these advances. For instance, Microsoft’s FarmBeats platform increased crop yields by 20% and reduced water use by 30% in U.S. and Indian trials (Microsoft Research, 2023). Retrieval-augmented advisory chatbots deployed in Kenya and India improved farmers’ decision-making speed and confidence in pest and weather management. A 2024 FAO pilot of digital sandboxes for pest control in Kenyan farms recorded an 18% reduction in pesticide use, demonstrating sustainable applications of GenAI.
RAI Challenges.
@@
Responsible adoption of GenAI in agriculture faces several interrelated challenges. These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions (Mamo et al., 2023b; Stojanovic et al., 2021), farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms (Bronson, 2019; Tzachor et al., 2022), bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers (Sapkota et al., 2024; Rotz et al., 2019), overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom (Bronson, 2019; Tzachor et al., 2022), IoT security vulnerabilities that expose connected farm systems to cyberattacks (Mahdavinejad et al., 2018; Khalid et al., 2023), and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; UNESCO, 2021). Empirical use cases underscore these risks. For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment. Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019). IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
RAI Strategies.
@@
Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture. Geo-aware retrieval and annotation techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks (Mamo et al., 2023a). Data governance contracts are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets (Tzachor et al., 2022; Bronson, 2019). Embedding human-in-the-loop oversight into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation (Stojanovic et al., 2021; He et al., 2023b). Privacy-preserving IoT integration, including differential privacy and federated learning, protects sensitive farm data across institutions (Kaissis et al., 2021; Khalid et al., 2023). Finally, alignment with sustainability regulations such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards (UNESCO, 2021; Tzachor et al., 2022). Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
@@ 5.6. Defense
GenAI is increasingly influencing defense through simulation, sensing, and decision support (U.S. Department of Defense, 2021; NATO, 2024). Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development (NATO Allied Command Transformation, 2024; Johns Hopkins Applied Physics Laboratory, 2025), threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness (Wortzel et al., 2025; National Institute of Standards and Technology, 2023), autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments (Defense Advanced Research Projects Agency, 2018, 2024), and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (U.S. Department of Defense, 2024; U.S. Department of Defense, Chief Digital and AI Office, 2024). Recent deployments provide evidence of these advances. For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018; Liu et al., [n. d.]). The U.S. Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (U.S. Army, 2024; News, 2024). NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis (NATO Allied Command Transformation, 2024, 2025; Huang et al., 2025a).
RAI Challenges.
@@
Responsible adoption of GenAI in defense introduces ethical operational and legal concerns. These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025), opacity of decision making that limits auditability and traceability in high-stakes contexts (U.S. Department of Defense, 2021; National Institute of Standards and Technology, 2023), adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning (National Institute of Standards and Technology, 2025a, 2024), escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk (Horowitz et al., 2024; Stockholm International Peace Research Institute, 2024), ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement (NATO, 2024; International Committee of the Red Cross, 2021), research security concerns over dual-use model capabilities and insider threat (U.S. Department of Defense, 2025; Corporation, 2024), protection of critical infrastructure as AI is integrated into command networks and sensors (National Institute of Standards and Technology, 2023; U.S. Department of Defense, 2025), and facial recognition deployment where demographic performance differences raise risks in security operations (Grother et al., 2019, 2022). Empirical use cases underscore these risks. For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024). Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defenses if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025). Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
RAI Strategies.
@@
Multiple strategies are being advanced to ensure responsible deployment. Human in the loop oversight is prioritized for mission-critical tasks so that trained operators validate model outputs before action (U.S. Department of Defense, 2021; NATO, 2024). Meaningful human control is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025). Explainable AI for defense transparency adopts traceability and interpretability requirements to support audit and after-action review (U.S. Department of Defense, 2021; National Institute of Standards and Technology, 2023). Red teaming and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding (Corporation, 2024; National Institute of Standards and Technology, 2025a). International cooperation aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024). Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S. Department of Defense, 2024; Defense Innovation Unit, 2020).
@@ 6.1. Synthesis and interpretation
Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. The distribution is uneven. Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention. This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2024–2025, which increases the need for evaluation that captures tool use and context.
6.2. Implications for research

### v6

@@ 5.5. Agriculture
GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation. Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage, crop yield prediction models that integrate weather, soil, and satellite data, pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy, virtual agronomy advisors that deliver context-sensitive recommendations to farmers, agricultural robotics for automated harvesting and monitoring, crop breeding simulations that accelerate trait discovery, and weather forecasting for climate adaptation (Costa et al., 2023). While platforms like Microsoft’s FarmBeats (Microsoft Research, 2023) initially achieved significant yield and water-saving gains through IoT and predictive ML , GenAI is now being integrated to augment these systems through generative weather simulation for risk modeling and as LLM-based natural language interfaces . These tools leverage GenAI to translate complex sensor data into actionable, localized advice.
ResGenAI challenges
@@
Responsible adoption of GenAI in agriculture faces several interrelated challenges. These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions, farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms, bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers, overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom, IoT security vulnerabilities that expose connected farm systems to cyberattacks, and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; Sapkota et al., 2024). Empirical use cases underscore these risks. For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment. Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019). IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
Evaluation context
Agriculture risks primarily concern C4 (hallucinated agronomic advice), C6 (farm data ownership and privacy), and C7 (system-level failure in IoT-integrated advisory pipelines). No benchmark in our assessment provides geo-aware hallucination testing, and C7 remains untested for sensor-to-advice chains. Priority KPIs are $E_{h}$ (KPI 9) operationalized as harmful agronomic recommendations per $N$ advisory sessions, $P$ (KPI 2) for data governance compliance, and $E_{\text{kWh}}$ (KPI 12) for sustainability reporting—particularly relevant given the energy constraints of low-resource deployment contexts.
RAI strategies
@@
Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture. Geo-aware retrieval and annotation (C4, KPI 9) techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks. Data governance contracts (C6, KPI 2) are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets. Embedding human-in-the-loop oversight (C7, KPI 10) into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation. Privacy-preserving IoT integration (C6, KPI 2), including differential privacy and federated learning, protects sensitive farm data across institutions. Finally, alignment with sustainability regulations (C10, KPI 12) such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards. Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
@@ 5.6. Defense
GenAI is increasingly influencing defense through simulation, sensing, and decision support. Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development, threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness, autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments, and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (News, 2024). Recent deployments provide evidence of these advances. For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018). The U.S. Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (News, 2024). NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis.
RAI challenges
@@
Responsible adoption of GenAI in defense introduces ethical operational and legal concerns. These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved, opacity of decision making that limits auditability and traceability in high-stakes contexts, adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning, escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk, ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement, research security concerns over dual-use model capabilities and insider threat, protection of critical infrastructure as AI is integrated into command networks and sensors, and facial recognition deployment where demographic performance differences raise risks in security operations. Empirical use cases underscore these risks. For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024). Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defences if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025). Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
Evaluation context
Defense risks concentrate on C3 (adversarial robustness), C7 (system-level failure in multi-agent autonomy), C8 (red-team realism), and C10 (governance alignment with IHL and NATO principles). C8 is the most critical gap: only 2 of 21 benchmarks (Rainbow Teaming and ALERT) score above zero for adaptive adversary simulation, yet defense applications face the most sophisticated threat actors. C7 is entirely absent for multi-agent autonomous systems. Priority KPIs are robustness $R$ (KPI 6) under military-grade adversarial scenarios, $E_{h}$ (KPI 9) for consequential targeting or escalation errors, and audit frequency $F_{a}$ (KPI 10) aligned with NATO after-action review cadences.
RAI strategies
@@
Multiple strategies are being advanced to ensure responsible deployment. Human in the loop oversight (C7, C8, KPI 10) is prioritized for mission-critical tasks so that trained operators validate model outputs before action. Meaningful human control (C10) is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law. Explainable AI for defense transparency (C9, KPIs 7–8) adopts traceability and interpretability requirements to support audit and after-action review. Red teaming (C3, C8, KPI 6) and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding. International cooperation (C10) aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024). Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S. Department of Defense, 2024; Defense Innovation Unit, 2020).
5.7. Cross-Applications Synthesis of GenAI Maturity
To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Table 9. Beyond risk mechanisms and regulatory drivers, the table now maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs, making explicit which evaluation evidence is available and where portable audit artifacts are missing.
5.7.1. Stakeholder-Metric Alignment: Operationalizing the Framework
To ensure the sectoral insights in Section 5 translate into actionable governance, we map the relationship between stakeholder roles and our evaluation infrastructure. Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor. In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 4 to guide iterative alignment during the fine tuning phase. Conversely, Compliance Auditors and Regulators prioritize the statistical parity (SPD/DI) KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and, NIST). Finally, Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the High-Stakes Error Rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of human-in-the-loop oversight. By bridging domain-specific mechanisms with normalized technical metrics, our framework provides the ”chain of trust” required to move ResGenAI from theoretical safety to verifiable institutional accountability.
Table 9. Synthesis Matrix: Comparative Analysis of RAI Maturity, Regulatory Pressures, Rubric Criteria Gaps, and Priority KPIs Across Domains.
Sector Dominant Risk Mechanism Technical Control Maturity Primary Regulatory Driver Primary C1 to C10 Gaps Priority KPIs
Healthcare Medical hallucination (Lyu et al., 2025); Diagnostic bias (Gonzalez et al., 2022) Medium: RAG (Wu et al., 2024), HITL, Diff. Privacy (Kaissis et al., 2021) FDA Guidance (U.S. Food and Drug Administration, 2024); HIPAA (Shiraz and Lee, 2019) C4: no clinical hallucination benchmark; C6: no HIPAA specific privacy suite; C7: 4/21 cover system failure $E_{h}$ (9), $P$ (2), SPD/DI (3 to 4), $\Delta_{\text{acc}}$ (5)
Finance Lending discrimination (Mehrabi et al., 2021b); Market manipulation (Hendry et al., 2022) High: Fairness audits (Pfohl et al., 2022); Multi party comp. (Khalid et al., 2023) SEC (U.S. Securities and Exchange Commission, 2021); OCC (Office of the Comptroller of the Currency, 2020) C1: no lending discrimination benchmark; C3: no financial adversarial attack suite SPD/DI (3 to 4), $R$ (6), $F_{a}$ (10)
Education Student privacy (Livingstone et al., 2021); Skills evasion (Pavlou et al., 2023) Low: Watermarking (Jiang et al., 2023); Fed. learning (Yang, 2021) UNESCO AI Rubric (UNESCO, 2021b); State Edu. Laws (Utah State Board of Education, 2022) C6: no minors data protocol; C7: skills evasion untested; KPI 11 lacks standard $P$ (2), $E_{h}$ (9), $U$ (11)
[…]
Agriculture Ownership conflicts (Bronson, 2019); Global South exclusion (Tzachor et al., 2022) Low: Geo aware RAG (Ma et al., 2023); Data contracts (Bronson, 2019) FAO (Costa et al., 2023); EU CAP (Costa et al., 2023) C4: no geo aware hallucination test; C7: IoT advisory pipelines untested $E_{h}$ (9), $P$ (2), $E_{\text{kwh}}$ (12)
Defense Autonomous lethality (International Committee of the Red Cross, 2021); Adversarial spoofing (National Institute of Standards and Technology, 2025a) High: Red teaming (Corporation, 2024); Meaningful HITL (International Committee of the Red Cross, 2021) NATO Principles (NATO, 2024); IHL (United Nations Institute for Disarmament Research, 2025) C8: only 2/21 benchmarks score >>0; C7: absent for multi agent autonomy $R$ (6), $E_{h}$ (9), $F_{a}$ (10)
Legal & Justice Recidivism prediction bias (Skeem and Lowenkamp, 2016); Automated decision opacity (Chesterman, 2021) Low to Med.: Explainability (LIME/SHAP) (Microsoft Research, [n. d.]); Bias audits (Center for Data Science and Public Policy, [n. d.]) EU AI Act Art. 6 (Edwards, 2021); U.S. DOJ Guidelines (Podgor, 2003) C1: no recidivism specific fairness benchmark; C2: legal advice toxicity untested; C10: judicial HITL criteria absent SPD/DI (3 to 4), $E_{h}$ (9), $R$ (6), $F_{a}$ (10)
Manufacturing Predictive maintenance failures (Hashemian, 2010); Adversarial sensor spoofing (Cao et al., 2019) Medium: Digital twins (Mihai et al., 2022); Anomaly detection (Chandola et al., 2009) IEC 62443 (ISO/IEC, 2021b); EU Machinery Reg. (European Union, 2024) C3: no industrial adversarial robustness suite; C4: no sensor drift benchmark; C7: multi robot coordination untested $E_{h}$ (9), $R$ (6), $E_{\text{kWh}}$ (12), $F_{a}$ (10)
@@ 6.1. Synthesis and Interpretation
Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention. This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2022–2025, which increases the need for evaluation that captures tool use and context.
Closing the accountability chain
To fully answer our primary research question, Who is Responsible?, our synthesis posits that the user must be transitioned from a passive consumer to a systemic guardrail. The current imbalance in GenAI governance, where 74% of organizations use GenAI but only 26% have a formal strategy, often results in the end-user inheriting unmitigated technical risks like PII leakage and hallucination. By triangulating the sectoral failures in Section 5 with the digital literacy gaps identified in Section 3.2.2, we propose a Symmetric Responsibility Model. In this model, developers are accountable for Transparency and Alignment, while users are accountable for Operational Integrity and Dissemination. Digital literacy, therefore, is not merely a skill but a core technical control that must be assessed alongside model-side benchmarks in Section 4.2. Without this user centric governance, even perfectly aligned models remain vulnerable to the last mile failure of human misinterpretation or malicious re contextualization. Table 10 operationalizes this model by mapping specific failure modes to the reciprocal duties of each stakeholder.
Table 10. The Symmetric Responsibility Matrix: Mapping Failure Modes to Stakeholder Duties
Failure Mode Developer Duty (Engineering Controls) User Duty (Operational Integrity)
Hallucination & Misinfo Implement RAG frameworks and calibrated uncertainty reporting (KPI 9) (Wu et al., 2024). Verify factual claims via primary sources and cross modal grounding checks.
PII & Data Leakage Deploy Differential Privacy and automated scrubbing during inference (KPI 2) (Nasr et al., 2023). Practice prompt hygiene by omitting sensitive or proprietary data from inputs.
Adversarial Jailbreaking Conduct continuous red teaming and implement robust system guardrails (C8) (Samvelyan et al., 2024). Adhere to acceptable use policies and report emergent safety bypasses to developers.
Representational Bias Perform fairness audits using intersectional metrics like SPD and DI (KPIs 3 to 4) (Kun, 2015). Critically appraise outputs for stereotyping and contextualize results within local norms.
Over reliance (Automation Bias) Provide model confidence scores and explicit explanation gap warnings (KPI 8) (Strathern, 1997a). Maintain human in the loop (HITL) oversight and preserve independent domain expertise.
Industry Safety Artifacts and Advanced Interpretability.
While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly defined by safety documentation from frontier labs and emerging mechanistic interpretability techniques, as synthesized in Figure 11. Frontier labs have established documentation protocols as de facto benchmarks for governance facing evidence, directly mapping to our C1 to C10 rubric. OpenAI system cards for GPT 4 and GPT 4o provide granular audits for sycophancy and PII leakage, demonstrating high maturity applications of metric validity through traceable safety refusal evidence. Anthropic utilizes Constitutional AI to automate alignment via RLAIF, outlining tiered safeguards that address malicious actor realism. Meta Llama 3 documentation introduces standardized protocols for multilingual red teaming, vital for addressing cross cultural ethical gaps. Google DeepMind technical reports provide empirical evidence for bias and accuracy trade offs, detailing the overcorrection mechanisms that led to observed representational failures. This industry evidence is complemented by the shift toward mechanistic interpretability, which maps internal neural activations to semantic concepts to offer a more rigorous safety check than behavior level prompts alone. Prominent techniques like Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling model steering to suppress harmful features during inference. Integrating these mechanistic audits into the RAI lifecycle addresses the black box challenge by enabling empirical verification of internal safety alignment before deployment.
Figure 11. Integration of industry safety documentation and advanced mechanistic interpretability for governance facing evidence.
ResGenAI sustainability via lifecycle assessment
Sustainability in ResGenAI must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into our governance landscape (Figure 8) and synthesized in Figure 12. This necessitates formalizing carbon reporting and energy benchmarks across the development cycle, including pretraining, fine tuning, and inference. Following our proposed lifecycle KPIs (Section 4.2.4), we advocate for reporting energy consumption and carbon equivalent emissions normalized by throughput, such as per 1K tokens generated. Seminal research indicates that training large scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, highlighting the need for transparent tracking tools and hardware aware reporting. A sustainable AI ecosystem requires a dual focus on optimizing the computational efficiency of models and leveraging generative insights to advance ecological goals, ensuring that technical innovation supports long term environmental stability.
Figure 12. Quantifying Responsible GenAI Sustainability via Lifecycle Assessment (LCA) in ResGenAI Governance. This framework formalizes carbon reporting and energy benchmarks across the pretraining, fine tuning, and inference stages, enabling a transition to rigorous quantification and supporting a sustainable AI ecosystem.
6.2. Lessons from Industry Deployment Incidents
While our benchmark analysis identifies risk-coverage gaps analytically, real-world deployment incidents from major AI developers offer empirical grounding. We synthesize prominent cases below, mapped to our rubric criteria, to demonstrate how the gaps in Tables 36 manifest in production systems.
Alignment and safety failures (C2, C3, C8) Microsoft Bing Chat integration of GPT-4 in early 2023 produced adversarial persona adoption and emotional manipulation during extended sessions (Roose, 2023), behaviors that emerged through multi-turn prompting, precisely the adaptive adversary pattern captured by C8, for which only 2 of 21 benchmarks scored above zero. OpenAI GPT-4o system card documented sycophantic and manipulative voice-mode outputs 444developers.openai.com/api/docs/models/gpt-4o, illustrating subtler alignment failures beyond toxicity classifiers. Google’s Gemini image generator was suspended in February 2024 after producing historically inaccurate racially diverse depictions of historical figures555blog.google/products/gemini/gemini-image-generation-issue/, demonstrating how overcorrection in bias mitigation (C1) can itself produce fairness failures, an interaction that benchmarks evaluate in isolation.
Hallucination in deployed systems (C4, C9) In Mata v. Avianca (2023), attorneys submitted a brief with six fabricated ChatGPT-generated case citations, resulting in judicial sanctions666caselaw.findlaw.com/court/us-dis-crt-sd-new-yor/2335142.html. These cases underscore that only 7 of 21 benchmarks address misinformation with documented constructs, and that the high-stakes error rate KPI ($E_{h}$, Section 4.2.4) must be operationalized in deployment settings.
Privacy and data leakage (C6) Samsung banned internal ChatGPT use in 2023 after employees inadvertently leaked proprietary source code through conversations (Samsung Electronics, 2023), a usage-path leakage vector that no current privacy benchmark assesses. Nasr et al. (Nasr et al., 2023) demonstrated scalable extraction of memorized training data, including PII, from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
Deepfake and media integrity (C5) A Hong Kong finance worker transferred approximately \$25 million after a video call with deepfake recreations of company executives (Chen, 2024), and AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024). Only SHIELD provides explicit C5 evaluation, and neither incident type, social engineering or democratic interference, falls within its scope.
AI Act enforcement (C10) The Act’s GPAI provisions, effective August 2025, require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024). The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 6. Early compliance experience confirms our finding that organizations struggle to assemble portable evidence across governance obligations, as no single benchmark suite covers more than half of the rubric at the “strong” level.
6.3. Trade-offs between Performance Measures
The Fairness-Privacy Paradox: Reconciling Representativeness with the Right to Anonymity
A fundamental tension emerging from our synthesis is the operational conflict between the requirement for Representativeness and the rigid Privacy mandates of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). While mitigating algorithmic bias (Section 4.2.1) necessitates the collection and processing of high-fidelity, demographically diverse datasets to ensure equitable model performance across protected classes, privacy-preserving standards often mandate data minimization and the obfuscation of sensitive attributes. This creates a structural “Fairness-Privacy Paradox”: without access to granular, protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional bias audits (Kun, 2015). Consequently, the legal push for anonymity under GDPR can inadvertently mask the very disparities that ResGenAI frameworks seek to eliminate. We posit that resolving this tension requires a transition from absolute data minimization to the adoption of privacy-enhancing technologies (PETs) such as Differential Privacy and Federated Learning that facilitate rigorous fairness auditing while maintaining individual-level anonymity .
Critical Appraisal of Technical Controls: Synthesis and Reflection
After looking closely at how these systems work in the real world, it’s clear there is a significant gap between what we hope these tools can do and how they actually perform when things get messy. The Explanation-Reliability Gap shows us that while we have many ways to explain model decisions (Dwivedi et al., 2023), these often create stories that sound good but don’t actually match what the model is doing under the hood (Strathern, 1997a). We have found that for big models, these explanations are mostly helpful for developers to find bugs, rather than giving regular people or regulators the true transparency they need.
The Bias-Overcorrection Paradox is another hard lesson we’ve learned from cases like the Gemini incident. When we try too hard to fix fairness using simple keywords, we often end up breaking the model’s ability to be factually accurate or culturally aware. This tells us that fairness isn’t just a number to hit (Kun, 2015), but a deep challenge that needs to change based on where the model is being used. Finally, The Shallow Guardrail Problem reminds us that models can look perfectly safe on paper (Zhang et al., 2024) while still being easy to trick. We’ve seen that models which pass every standard toxicity test can still be pushed into strange or harmful behaviour through basic multi-turn conversations (Roose, 2023). The reality is that our current ”guardrails” often just polish the output without fixing the underlying issues, leaving them vulnerable to the automated attacks we see growing today (Samvelyan et al., 2024).
6.4. Analytical Triangulation: Bridging the Evaluation Gap
The persistent safety failures in production systems are not merely isolated technical oversights; they are symptoms of a systemic “Evaluation Gap” where current benchmarks fail to model the adaptive, iterative pressure applied by real-world adversaries. This section triangulates the theoretical risk taxonomy in Section 3 with the empirical deployment failures documented in Section 6.2 through the lens of our comparative assessment (Section 4.2).
Our assessment in Table 4 reveals a stark deficit in evaluative maturity regarding Malicious Actor Realism (Criterion C8), where only 2 out of 21 surveyed benchmarks: Rainbow Teaming and ALERT provide strong coverage for adaptive pressure. This misalignment directly correlates with the high-profile incidents synthesized in Section 6.2, such as the adversarial persona adoption observed in Bing Chat (Roose, 2023) and the scalable data extraction vulnerabilities reported in ChatGPT (Nasr et al., 2023). These failures underscore that prevailing evaluation suites remain largely static and task-local, leaving them unequipped to measure the non-deterministic tactics encountered in the wild.
Furthermore, the representational failures observed in the Gemini overcorrection incident illustrate a critical breakdown in Governance Alignment (C10), where technical alignment efforts are decoupled from the broader sociotechnical context of deployment. By triangulating these findings, we posit that a ResGenAI operating model must transition from one-shot static probes toward continuous, adaptive evaluation frameworks that mirror the malicious-actor realism and distribution shifts inherent in generative ecosystems.
6.5. Limitations of the Study
Like any survey-based study, this work has several limitations. First, the review follows a pre-specified protocol and focuses primarily on the period from November 2022 to December 2025, with selective inclusion of foundational works for definitional clarity. While this post-ChatGPT window captures the most rapid phase of GenAI deployment, it may exclude emerging studies published after the search cutoff or earlier work that did not explicitly use contemporary terminology. Second, our synthesis relies on publicly available papers, dataset cards, benchmark documentation, and governance artifacts. As a result, proprietary evaluations, internal red-teaming practices, or unpublished safety mitigations may be under-represented. Documentation quality also varies across benchmarks and frameworks, which may affect the completeness of criterion-level scoring.
Third, rubric weights are assigned equally by default, and criterion coverage is determined through structured reading of sources followed by discussion and adjudication in cases of disagreement. Although we conducted sensitivity analyses to assess the robustness of equal-weight assumptions, alternative weighting schemes reflecting sector-specific risk priorities could yield different relative rankings. Fourth, the scoring procedure captures breadth of coverage rather than depth or real-world effectiveness. A benchmark receiving a high rubric score indicates multi-risk awareness and governance alignment, but does not guarantee empirical robustness under adversarial deployment conditions. In this sense, the rubric reflects documented capability rather than independently verified performance.
Fifth, the review synthesizes heterogeneous evidence types, including empirical evaluations, conceptual frameworks, and regulatory analyses. This diversity strengthens thematic coverage but precludes formal meta-analysis or quantitative effect-size aggregation. Our conclusions therefore emphasize structural patterns rather than causal claims. Sixth, governance alignment is assessed at the level of plausible evidentiary support rather than legal compliance. Mapping benchmarks to instruments such as the EU AI Act or NIST AI RMF indicates potential contribution to audit evidence, but does not constitute certification or conformity assessment.
Finally, the rapid evolution of multimodal and agentic systems means that new failure modes may emerge after publication. Continuous updates to evaluation suites, governance frameworks, and deployment practices may shift the relative landscape described here. Replication of our screening and scoring is facilitated by the availability of documented sources, explicit inclusion criteria, and the published rubric; however, interpretive judgment remains inherent in structured qualitative synthesis.
6.6. Research directions